OpenStudy (anonymous):

hi.. please I need this: indefinite integration of(x^2 + 2x + 1 / x^2 + 1)

OpenStudy (anonymous):

indefinite integral of (x^2+2x+1)/(x^2+1) indefinite integral of (x+1)(x+1)/(x+1)(x-1) indefinite integral of (x+1)/(x-1) I feel bad, but after this I can't quite finish the integration

OpenStudy (turingtest):

long division\[\frac{x^2+2x+1}{x^2+1}=1+\frac{2x}{x^2+1}\]now integrate

OpenStudy (anonymous):

thanks